# QCResUNet: Joint Subject-level and Voxel-level Segmentation Quality Prediction

### v1

@@ 2.3.1 Network Design
Encoder: For the purpose of predicting subject-level DSC, we adopted a ResNet-34 (He et al., 2016) architecture as part of the encoding path of our network, which can capture semantically rich features that are important for accurately characterizing segmentation quality. While retaining the main structure of the 2D ResNet-34 in He et al. (2016), we made the following modifications to account for the 3D nature of the input data. First, all 2D convolutional and pooling layers were replaced by 3D counterparts (see Figure 2(b)). Second, we replaced all batch normalization blocks (Ioffe and Szegedy, 2015) with instance normalization blocks (Ulyanov et al., 2016) to cater to the small batch size during 3D model training. Third, to prevent overfitting, spatial dropout (Tompson et al., 2015) was added to each residual block with a probability of 0.3 to randomly zero out channels in the feature map (see Figure 2(b)).
@@
Decoder: To estimate a segmentation error map with high spatial resolution and accurate localization information, we designed a decoder architecture. The decoder takes the low-resolution contextual features of segmentation quality that were extracted by the encoder and transfers them to a high-resolution multiclass SEM. This was accomplished by first upsampling the input feature map by a factor of two using nearest neighbor interpolation, followed by a $1\times 1\times 1$ convolutional layer. Next, we concatenated the upsampled feature maps with the corresponding encoder level’s features using skip-connections to facilitate information flow from the encoder to the decoder. This was followed by two convolutional blocks that reduced the number of feature maps back to the original value before concatenation. Each convolutional block comprised a $3\times 3\times 3$ convolutional layer, followed by an instance normalization layer and a Leaky Rectified Linear Unit (LeakyReLU) activation function (Maas et al., 2013) (see Figure 2(c)).
@@
Importantly, we used the middle-level semantics in the intermediate block (Block 3 in Figure 2(a)) of the ResNet encoder as the input to the decoder. The rationale behind this was that the last block (Block 4 in Figure 2(a)) of the ResNet encoder contains features that are specific to the DSC prediction task. In contrast, the middle-level features are likely to constitute a more universal semantic representation that characterizes the segmentation error (Ahn et al., 2019).
@@
Attention-based SEM aggregation: To better predict the segmentation error map corresponding to different tissue classes (e.g., [$\text{SEM}_{\text{WT}}$, $\text{SEM}_{\text{TC}}$, $\text{SEM}_{\text{ET}}$] in the brain tumor segmentation QC task), we propose an attention-based SEM aggregation mechanism. This is because each tissue class in the query segmentation mask ($S_{query}$) contributes differently to the predicted SEM. Specifically, we leveraged the efficient channel attention (ECA) (Wang et al., 2020a) to model the correlation between the features in the last layer of the decoder and the one-hot encoded query segmentation mask. A $1\times 1\times 1$ convolution layer was then applied to the output of the ECA module to predict tissue-level SEMs (Figure 2(a)). Specifically, the channel attention was computed by feeding the pooled input feature (after the global average pooling layer) to a 1D convolutional layer followed by a Sigmoid activation (Figure 2(d)).
@@ 2.3.3 Multi-task learning objective
where $N$ is the total number of samples in the training dataset, and $n$ is indexing each sample. The MAE loss ($\mathcal{L}_{MAE}$) quantifies the dissimilarities between the ground truth DSC/NSD (DSCgt/NSDgt) and the predicted DSC/NSD (DSCpred/NSDpred). Compared to the mean squared error loss commonly used for regression tasks, the MAE loss has been proved to be more sensitive to outliers (Qi et al., 2020). This enhances the robustness of QCResUNet to outliers.
The voxel-level segmentation error prediction task was trained by optimizing both the DSC and CE losses. Though there are many variants of the dice loss, we opted for the one proposed in (Drozdzal et al., 2016) due to its wide success in a variety of medical imaging segmentation tasks:
@@ 3.1 Brain tumor MRI segmentation QC task
We performed hyperparameter tuning to determine two critical hyperparameters in the proposed method: the initial learning rate and the loss balance parameter $\lambda$. Hyperparameter tuning was carried out by performing a Bandit-based Bayesian optimization on the training and validation dataset implemented using the Raytune framework (Liaw et al., 2018). Specifically, we sampled the learning rate from a range of $[1\times 10^{-5},2\times 10^{-4}]$ following a log-uniform distribution. Additionally, the loss balance parameter $\lambda$ was chosen from the following set of values $\{0.1,1,2\}$. The optimal balance parameter $\lambda$ for QCResUNet was determined to be 1. However, we did not observe a significant difference when varying $\lambda$ from 0.1 to 1 and 2 (see Figure S1). Although the optimal learning rate varied from model to model (refer to Table 2), it generally fell within the scale of $1\times 10^{-4}$ as we used a relatively small batch size of 4. We kindly direct the readers to C for the detailed learning curve in hyperparameter tuning using RayTune.
@@ 3.3 Baseline methods
To evaluate the effectiveness of the proposed model, we compared its performance with five baseline models: (i) RCA method (Robinson et al., 2017); (ii) uncertainty estimation (UE) based method (Jungo et al., 2020); and three regression-based methods: (iii) a UNet model (Ronneberger et al., 2015); (iv) a ResNet-34 model (He et al., 2016); and (v) a ResNet-50 (He et al., 2016) model (Robinson et al., 2018).
@@
RCA for brain tumor segmentation QC was implemented using as segmentation method the DeepMedic framework, which was adapted following the protocol described by Valindria et al. (2017). Specifically, we reduced the number of feature maps in each layer by one third compared to the default setting of DeepMedic. We also reduced the feature maps in the last fully connected layer from 150 to 45. We opted for this approach instead of the atlas-based segmentation approach because establishing spatial correspondences between query and reference images in tumorous areas can be challenging due to the high spatial and phenotypic heterogeneity of brain tumors. In contrast, RCA for cardiac MRI segmentation QC was implemented using the atlas-based segmentation method by following the protocols in (Robinson et al., 2019). This is because registration correspondences between query and reference are easier to establish in healthy autonomy. For the selection of the reference data set, we randomly selected 100 samples from the training dataset following a previously well-validated protocol (Robinson et al., 2019).
@@
The implementation of the UE-based method followed the protocol in (Jungo et al., 2020), where the uncertainty map was calibrated by evaluating the uncertainty-error (UE) overlap. The UE overlap measures the overlap between the binarized uncertainty map and the corresponding SEM (see B.1 for details). We chose the UE calibration proposed by Jungo et al. (2020) instead of other calibration methods (Wen et al., 2020; Ashukha et al., 2020; Mehta et al., 2022) because it produces a binary mask by thresholding the uncertainty map. We can directly compare the resulting binary mask to the SEM generated by the proposed QCResUNet. The subject-level DSC was computed by applying a random forest with 102 extracted radiomics features (see B.2).
@@
Figure 3: Scatter plots of the ground-truth (x-axis) and the predicted DSC (y-axis) for the proposed method and all the baseline methods in the internal BraTS testing, external BraTS-SSA, WUSM datasets, and ACDC internal testing (rows). Results for (a) RCA; (b) UE-based; (c) UNet; (d) ResNet-34; (e)ResNet-50; and (f) QCResUNet are reported in different columns. The proposed method generalized well to external datasets and consistently showed superior performance compared to all baseline methods. The UE-based method showed the worst performance compared to all other methods.
@@
Figure 4: Scatter plots of the ground-truth (x-axis) and the predicted NSD (y-axis) for the proposed method and all the baseline methods in the internal BraTS testing, external BraTS-SSA, WUSM datasets, and ACDC internal testing (rows). Results for (a) RCA; (b) UE-based; (c) UNet; (d) ResNet-34; (e)ResNet-50; and (f) QCResUNet are reported in different columns. The proposed method generalized well to external datasets and consistently showed superior performance compared to all baseline methods. The UE-based method showed the worst performance compared to all other methods.
@@ 3.5 Implementation details
Training procedures were the same for both tasks by using segmentation results were produced by nnUNet, nnFormer, and SegGen. Training was carried out using Adam optimizer (Kingma and Ba, 2014) with a batch size of 4 for 100 epochs. Training was started with an optimal initial learning rate determined by the hyperparameter search and was exponentially decayed by a factor of 0.9 for each epoch until it reached $1\times 10^{-6}$. We applied an $L_{2}$ weight decay of $1\times 10^{-4}$ in all trainable layers. To prevent overfitting, various data augmentations, including random rotation, scaling, mirroring, Gaussian noise, and gamma intensity correction, were applied during training. All model training was performed on NVIDIA Tesla A100 GPUs. By default, we used a mix of single-precision (FP32) and half-precision (FP16) tensors during training to reduce training time and memory consumption. The proposed method was implemented using PyTorch v1.12.1. All the code and pre-trained models developed for this work will be available at https://github.com/sotiraslab/QCResUNet.
@@ 4.1.1 Brain tumor MRI segmentation QC task
Importantly, the proposed method outperformed all baseline methods in NSD and DSC prediction tasks. Compared to the three regression-based QC methods (i.e., UNet, ResNet-34, ResNet-50), QCResUNet improved the second-best method by 0.8% for NSD prediction and 1.5% for DSC prediction in terms of Pearson r values on the BraTS internal testing set. On the external BraTS-SSA and WUSM datasets, QCResUNet outperformed the second-best method by an average of 1.3% and 1.9% in terms of Pearson r, respectively. A paired t-test confirmed that this improvement was statistically significant compared to all three regression-based methods baseline (Table 3). The proposed approach exhibited more evenly distributed DSC prediction errors across different quality levels compared to all the baseline methods (refer to Figure 3 and Figure 4), demonstrating a smaller standard deviation in MAE (see Table 3).
@@
Figure 6: Examples showcasing the performance of the proposed method versus baseline methods on the brain tumor segmentation QC task: (a) high-quality segmentation error localization and (b) low-quality segmentation error localization. The color bar in the figure indicates the intensity of the uncertainty map (UMap). We observed that the proposed method showed better segmentation error localization than the uncertainty map. The error localization was better when dealing with low-quality query segmentation in contrast to higher-quality ones. This may be attributed to the fact that detecting few errors at the boundaries of high-quality segmentations is challenging. We kindly direct the readers to D for details on how the visualization in this figure was obtained.
Figure 7: Examples of the class activation maps produced by Grad-CAM for different methods. The CAMs of UNet, ResNet-34, and ResNet-50 were generated based on the last convolutional feature maps. The CAMs of the proposed QCResUNet were generated from the last convolutional feature map of the ResNet encoder. We observed that the proposed QCResUNet showed superior localization performance in comparison to all other baselines in terms of the CAM. This may be attributed to the multi-task learning framework of the proposed method (see detailed discussion in Section 4.4).
The proposed method demonstrated a strong performance gain over the state-of-the-art RCA and UE-based QC methods (Table 3). The performance of the RCA and the UE-based method after hyper-parameter tuning was in line with previous works (Robinson et al., 2017; Jungo et al., 2020). In the internal BraTS testing set, the proposed method improved the average Pearson r of NSD predictions by 48.5% and DSC predictions by 22.1% compared to RCA and UE-based QC, respectively. A similar trend was observed on the external datasets, with the proposed method improving Pearson $r$ by an average of $50.9\%$ and $48.2\%$ compared to RCA and UE-based QC, respectively. Moreover, the proposed method achieved a significant reduction in the average MAE of predicting NSD and DSC compared to the RCA and UE-based methods for all results viz. internal testing results ($0.292$, $0.126$ vs. $0.064$), BraTS-SSA results ($0.281$, $0.157$ vs. $0.059$), and WUSM results ($0.275$, $0.179$ vs. $0.081$).
@@ 4.1.2 Cardiac MRI segmentation QC task
Similar to the brain tumor segmentation QC task, the proposed method achieved a good performance on the cardiac segmentation QC task (see Table 4). Specifically, the proposed method achieved average MAEs of 0.070 and 0.057 for NSD and DSC predictions, respectively. The predicted NSD and DSC demonstrated a strong correlation with the corresponding ground truth, with Pearson r values of 0.914 and 0.955, respectively. The proposed QCResUNet also outperformed all baseline methods in the cardiac segmentation QC task. QCResUNet improved the second-best regression-based method by 2.6% and 1.8% in Pearson r for NSD and DSC predictions, respectively. In addition, QCResUNet significantly outperformed RCA and UE-based QC methods by an average of 10.2% and 10.9%. Lastly, the proposed method achieved a significant reduction in MAEs compared to all baselines (Table 4), aligning well with the ground truth segmentation quality measures (see Figure 3 and Figure 4).

### v2

@@ 2.3.1 Network Design
Encoder: For the purpose of predicting subject-level DSC, we adopted a ResNet-34 [20] architecture as part of the encoding path of our network, which can capture semantically rich features that are important for accurately characterizing segmentation quality. While retaining the main structure of the 2D ResNet-34 in [20], we made the following modifications to account for the 3D nature of the input data. First, all 2D convolutional and pooling layers were replaced by 3D counterparts (see Figure 2(b)). Second, we replaced all batch normalization blocks [21] with instance normalization blocks [56] to cater to the small batch size during 3D model training. Third, to prevent overfitting, spatial dropout [54] was added to each residual block with a probability of 0.3 to randomly zero out channels in the feature map (see Figure 2(b)).
@@
Decoder: To estimate a segmentation error map with high spatial resolution and accurate localization information, we designed a decoder architecture. The decoder takes the low-resolution contextual features of segmentation quality that were extracted by the encoder and transfers them to a high-resolution multiclass SEM. This was accomplished by first upsampling the input feature map by a factor of two using nearest neighbor interpolation, followed by a $1\times 1\times 1$ convolutional layer. Next, we concatenated the upsampled feature maps with the corresponding encoder level’s features using skip-connections to facilitate information flow from the encoder to the decoder. This was followed by two convolutional blocks that reduced the number of feature maps back to the original value before concatenation. Each convolutional block comprised a $3\times 3\times 3$ convolutional layer, followed by an instance normalization layer and a Leaky Rectified Linear Unit (LeakyReLU) activation function [37] (see Figure 2(c)).
@@
Importantly, we used the middle-level semantics in the intermediate block (Block 3 in Figure 2(a)) of the ResNet encoder as the input to the decoder. The rationale behind this was that the last block (Block 4 in Figure 2(a)) of the ResNet encoder contains features that are specific to the DSC prediction task. In contrast, the middle-level features are likely to constitute a more universal semantic representation that characterizes the segmentation error [2].
@@
Attention-based SEM aggregation: To better predict the segmentation error map corresponding to different tissue classes (e.g., [$\text{SEM}_{\text{WT}}$, $\text{SEM}_{\text{TC}}$, $\text{SEM}_{\text{ET}}$] in the brain tumor segmentation QC task), we propose an attention-based SEM aggregation mechanism. This is because each tissue class in the query segmentation mask ($S_{query}$) contributes differently to the predicted SEM. Specifically, we leveraged the efficient channel attention (ECA) [60] to model the correlation between the features in the last layer of the decoder and the one-hot encoded query segmentation mask. A $1\times 1\times 1$ convolution layer was then applied to the output of the ECA module to predict tissue-level SEMs (Figure 2(a)). Specifically, the channel attention was computed by feeding the pooled input feature (after the global average pooling layer) to a 1D convolutional layer followed by a Sigmoid activation (Figure 2(d)).
@@ 2.3.3 Multi-task learning objective
where $N$ is the total number of samples in the training dataset, and $n$ is indexing each sample. The MAE loss ($\mathcal{L}_{MAE}$) quantifies the dissimilarities between the ground truth DSC/NSD (DSCgt/NSDgt) and the predicted DSC/NSD (DSCpred/NSDpred). Compared to the mean squared error loss commonly used for regression tasks, the MAE loss has been proved to be more sensitive to outliers [42]. This enhances the robustness of QCResUNet to outliers.
The voxel-level segmentation error prediction task was trained by optimizing both the DSC and CE losses. This is because combining CE and DSC loss has become mainstream for training models involving dense voxel/pixel-level predictions [23, 63, 12, 9, 36]. Though there are many variants of the dice loss, we opted for the one proposed in [14] due to its wide success in a variety of medical imaging segmentation tasks:
@@ 3.1 Brain tumor MRI segmentation QC task
We performed hyperparameter tuning to determine two critical hyperparameters in the proposed method: the initial learning rate and the loss balance parameter $\lambda$. Hyperparameter tuning was carried out by performing a Bandit-based Bayesian optimization on the training and validation dataset implemented using the Raytune framework [34]. Specifically, we sampled the learning rate from a range of $[1\times 10^{-5},2\times 10^{-4}]$ following a log-uniform distribution. Additionally, the loss balance parameter $\lambda$ was chosen from the following set of values $\{0.1,1,2\}$. The optimal balance parameter $\lambda$ for QCResUNet was determined to be 1. However, we did not observe a significant difference when varying $\lambda$ from 0.1 to 1 and 2 (see Figure S1). Although the optimal learning rate varied from model to model (refer to Table 3), it generally fell within the scale of $1\times 10^{-4}$ as we used a relatively small batch size of 4. We kindly direct the readers to C for the detailed learning curve in hyperparameter tuning using RayTune.
@@ 3.3 Baseline methods
To evaluate the effectiveness of the proposed model, we compared its performance with five baseline models: (i) RCA method [47]; (ii) uncertainty estimation (UE) based method [24]; and three regression-based methods: (iii) a UNet model [49]; (iv) a ResNet-34 model [20]; and (v) a ResNet-50 [20] model [45].
@@
RCA for brain tumor segmentation QC was implemented using as segmentation method the DeepMedic framework, which was adapted following the protocol described by [57]. Specifically, we reduced the number of feature maps in each layer by one third compared to the default setting of DeepMedic. We also reduced the feature maps in the last fully connected layer from 150 to 45. We opted for this approach instead of the atlas-based segmentation approach because establishing spatial correspondences between query and reference images in tumorous areas can be challenging due to the high spatial and phenotypic heterogeneity of brain tumors. In contrast, RCA for cardiac MRI segmentation QC was implemented using the atlas-based segmentation method by following the protocols in [46]. This is because registration correspondences between query and reference are easier to establish in healthy autonomy. For the selection of the reference data set, we randomly selected 100 samples from the training dataset following a previously well-validated protocol [46].
@@
The implementation of the UE-based method followed the protocol in [24], where the uncertainty map of each generated segmentation sample was obtained by applying the MCDropout [16] to the trained nnUNet, nnFormer and DeepMedic (see B). The uncertainty map was then calibrated by evaluating the uncertainty-error (UE) overlap. The UE overlap measures the overlap between the binarized uncertainty map and the corresponding SEM (see B.1 for details). We chose the UE calibration proposed by [24] instead of other calibration methods [62, 4, 39] because it produces a binary mask by thresholding the uncertainty map. We can directly compare the resulting binary mask to the SEM generated by the proposed QCResUNet. The subject-level DSC was computed by applying a random forest with 102 extracted radiomics features (see B.2).
@@
Fig. 3: Scatter plots of the ground-truth (x-axis) and the predicted DSC (y-axis) for the proposed method and all the baseline methods in the internal BraTS testing, external BraTS-SSA, WUSM datasets, and ACDC internal testing (rows). Results for (a) RCA; (b) UE-based; (c) UNet; (d) ResNet-34; (e)ResNet-50; and (f) QCResUNet are reported in different columns. The proposed method generalized well to external datasets and consistently showed superior performance compared to all baseline methods. The UE-based method showed the worst performance compared to all other methods.
@@
Fig. 4: Scatter plots of the ground-truth (x-axis) and the predicted NSD (y-axis) for the proposed method and all the baseline methods in the internal BraTS testing, external BraTS-SSA, WUSM datasets, and ACDC internal testing (rows). Results for (a) RCA; (b) UE-based; (c) UNet; (d) ResNet-34; (e)ResNet-50; and (f) QCResUNet are reported in different columns. The proposed method generalized well to external datasets and consistently showed superior performance compared to all baseline methods. The UE-based method showed the worst performance compared to all other methods.
@@ 3.5 Implementation details
Training procedures were the same for both tasks by using segmentation results were produced by nnUNet, nnFormer, and SegGen. Training was carried out using Adam optimizer [28] with a batch size of 4 for 100 epochs. Training was started with an optimal initial learning rate determined by the hyperparameter search and was exponentially decayed by a factor of 0.9 for each epoch until it reached $1\times 10^{-6}$. We applied an $L_{2}$ weight decay of $1\times 10^{-4}$ in all trainable layers. To prevent overfitting, various data augmentations, including random rotation, scaling, mirroring, Gaussian noise, and gamma intensity correction, were applied during training. All model training was performed on NVIDIA Tesla A100 GPUs. By default, we used a mix of single-precision (FP32) and half-precision (FP16) tensors during training to reduce training time and memory consumption. The proposed method was implemented using PyTorch v1.12.1.
@@ 4.1.1 Brain tumor MRI segmentation QC task
Importantly, the proposed method outperformed all baseline methods in NSD and DSC prediction tasks. Compared to the three regression-based QC methods (i.e., UNet, ResNet-34, ResNet-50), QCResUNet improved the second-best method by 0.8% for NSD prediction and 1.5% for DSC prediction in terms of Pearson r values on the BraTS internal testing set. On the external BraTS-SSA and WUSM datasets, QCResUNet outperformed the second-best method by an average of 1.3% and 1.9% in terms of Pearson r, respectively. A paired t-test confirmed that this improvement was statistically significant compared to all three regression-based methods baseline (Table 4). In addition, the proposed method outperformed multi-dimensional regression-based method [15] in subject-level DSC prediction by an average of 56.3%, 58.8%, and 180% on internal testing, external BraTS-SSA, and external WUSM datasets, respectively (Table 4). We hypothesize that the inferior performance of multi-dimensional regression-based methods, compared to other regression-based approaches, is attributed to the aggregation of 2D DSCs across slices without accounting for tumor sizes. As a consequence, each slice contributes equally to the aggregation of 2D DSCs. However, the variability in tumor sizes within a single 3D subject or across different subjects presents a significant challenge in accurately aggregating them for subject-level DSC prediction. The proposed approach exhibited more evenly distributed DSC prediction errors across different quality levels compared to all the baseline methods (refer to Figure 3 and Figure 4), demonstrating a smaller standard deviation in MAE (see Table 4). Furthermore, our QCResUNet offers computational efficiency comparable to ResNet50 and slightly surpasses that of UNet, especially in terms of FLOPs (see more computational benchmarking in F).
@@
The proposed method demonstrated a strong performance gain over the state-of-the-art RCA and UE-based QC methods (Table 4). The performance of the RCA and the UE-based method after hyper-parameter tuning was in line with previous works [47, 24]. In the internal BraTS testing set, the proposed method improved the average Pearson r of NSD predictions by 48.5% and DSC predictions by 22.1% compared to RCA and UE-based QC, respectively. A similar trend was observed on the external datasets, with the proposed method improving Pearson $r$ by an average of $50.9\%$ and $48.2\%$ compared to RCA and UE-based QC, respectively. Moreover, the proposed method achieved a significant reduction in the average MAE of predicting NSD and DSC compared to the RCA and UE-based methods for all results viz. internal testing results ($0.292$, $0.126$ vs. $0.064$), BraTS-SSA results ($0.281$, $0.157$ vs. $0.059$), and WUSM results ($0.275$, $0.179$ vs. $0.081$).
@@ 4.1.2 Cardiac MRI segmentation QC task
Similar to the brain tumor segmentation QC task, the proposed method achieved a good performance on the cardiac segmentation QC task (see Table 5). Specifically, the proposed method achieved average MAEs of 0.070 and 0.057 for NSD and DSC predictions, respectively. The predicted NSD and DSC demonstrated a strong correlation with the corresponding ground truth, with Pearson r values of 0.914 and 0.955, respectively. The proposed QCResUNet also outperformed all baseline methods in the cardiac segmentation QC task. QCResUNet improved the second-best regression-based method(i.e., ResNet50 [45]) by 2.6% and 1.8% in Pearson r for NSD and DSC predictions, respectively. In contrast to poor performance in the brain tumor segmentation QC task, the multi-dimensional regression-based method [15] performed reasonably well. Despite showing inferior performance compared to other regression-based methods, it outperformed RCA and UE-based methods (see Table 5).
In addition, QCResUNet significantly outperformed RCA and UE-based QC methods by an average of 10.2% and 10.9%. Lastly, the proposed method achieved a significant reduction in MAEs compared to all baselines (Table 5), aligning well with the ground truth segmentation quality measures (see Figure 3 and Figure 4).
Fig. 6: Examples showcasing the performance of the proposed method versus baseline methods on the brain tumor segmentation QC task: (a) high-quality segmentation error localization and (b) low-quality segmentation error localization. The color bar in the figure indicates the intensity of the uncertainty map (UMap). We observed that the proposed method showed better segmentation error localization than the uncertainty map. The error localization was better when dealing with low-quality query segmentation in contrast to higher-quality ones. This may be attributed to the fact that detecting few errors at the boundaries of high-quality segmentations is challenging. We kindly direct the readers to D for details on how the visualization in this figure was obtained.
Fig. 7: Examples of the class activation maps produced by Grad-CAM for different methods, with regions of interest (ROIs) highlighted in purple boxes. The CAMs of UNet, ResNet-34, and ResNet-50 were generated based on the last convolutional feature maps. The CAMs of the proposed QCResUNet were generated from the last convolutional feature map of the ResNet encoder. We observed that the proposed QCResUNet demonstrated improved localization performance compared to the other baselines in terms of the CAM. This may be attributed to the multi-task learning framework of the proposed method (see detailed discussion in Section 4.4).